# Magnetic field amplification and decay in cosmic string wakes

### v2

@@ Abstract
We do a detailed study on vortex formation in a magnetized plasma within the spacetime of a moving cosmic string using analytical and numerical methods. The conical spacetime around the cosmic string causes the frozen-in magnetic field to deform due to the fluid flow. We find that the overdensity in the wake region amplifies the magnetic field. This amplification depends on the direction and the lengthscale of the magnetic perturbations. Alfven’s theorem of flux conservation explains this result. However, our study also shows that the magnetic field can decay depending on the perturbation lengthscale, due to the breakdown of Alfven’s theorem at a certain lengthscale. This lengthscale is the gyroradius of the charged particles in the plasma. Our findings are significant for understanding magnetic reconnection in cosmic string wakes.
journal: Eur. Phys. J. C
1 Introduction
@@
Cosmic strings are topological defects that are formed from the symmetry-breaking phase transitions in the early universe. An interesting part of these defects is the nature of the spacetime around the string. It is known that the cosmic string metric is globally conical and has a deficit angle. Due to the nature of the metric, the geodesics of particles close to the cosmic string move inwards towards the cosmic string. This gives rise to a wake structure behind a moving cosmic string. This wake structure is well studied for an un-magnetized plasma vishniac ; sornborger ; stebbins . These wakes leave several signatures during the evolution of the universe. They can give rise to discontinuties in the Cosmic Microwave Background Radiation (CMBR)danos ; they may give rise to primordial magnetic fields vachaspati1 ; vachaspati2 ; sovan and they may also lead to density inhomogeneities which will affect the nucleosynthesis calculations layek1 ; layek ; sovan1 . Thus, understanding the structure of these wakes is important for identifying the signatures of cosmic strings in the current universe. The wake structure that has been studied in the literature is essentially the hydrodynamic wake structure. The presence of magnetic fields in the wake was generally attributed only to the superconducting cosmic strings rogozin . However, recently it has been shown that primordial magnetic fields can be also be generated in the wakes of Abelian Higgs strings sovan . So, it is possible that wakes of non-superconducting strings will also have a magnetic field. Particles moving through these magnetic fields will radiate synchrotron radiation. Such radiation can be detected by current detectors. We need to study the electron distribution, the magnetic field perturbation and many other parameters before we are able to obtain a distinct signature that points to a cosmic string. Our motivation in this study is to study in greater detail the flow of the magnetized fluid around a cosmic string and understand the important lengthscales associated with this flow.
@@
Recent studies has shown that multiple shocks can be generated in the cosmic string wake soumen . These shocks can collide and release a large amount of energy as radiation. Moreover, recently it has also been conjectured that the narrow wake structure of the cosmic strings can lead to magnetic reconnection dilip . Lengthscales play an important role in magnetic reconnection. One of the requirements of magnetic reconnection is the breakdown of Alfven’s theorem in the magnetized plasma. This can happen at small lengthscales. We aim to identify the lengthscale at which magnetic reconnection can occur in cosmic string wakes. This lenghscale is important as it determines the energy released due to magnetic reconnection.
@@
The fluid flow around the cosmic string is driven by the conical metric of the cosmic string. The magnetic field lines in the early universe are usually frozen into the plasma. Studies have shown that when the plasma undergoes a spherical compression, the magnetic field frozen in the plasma is also affected king . Since the flow around the cosmic string results in the fluid density increasing significantly in the wake region, we do a detailed study of the evolution of frozen-in magnetic fields around a cosmic string.
@@
In the literature, the magnetic field in the wake region is generated due to the presence of density inhomogeneities in the plasma. These density inhomogeneities come from the clustering of different particles around the cosmic string. The clustering of these particles mostly depend on their masses abhisek . A long and narrow wake is formed. In our numerical simulations, we have found that when the velocity of the cosmic string is low, the wake structure is diffuse. A well developed wake structure is obtained for higher Mach numbers. For simplicity, we consider an ideal fluid approximation without any dissipative forces. To generate some vorticity in the wake, we give a small velocity perturbation to the particle velocities in both the directions in the plane of the wake. The small vorticity generated in the plasma is conserved. Since there are no turbulent forces, we do not study the evolution of the vorticity with time. Our main aim is to study the evolution of the magnetic field over time in the absence of any turbulent forces.
We find that the field in the wake region can increase even without the presence of any external driving force or the presence of a dynamo mechanism. This is expected as the field lines are frozen in and so the conservation of magnetic flux density usually enhances the field when the fluid flow squeezes the flux lines. So, it is not surprising that the magnetic field is amplified due to the overdensity of the wake. However, we also find that in certain cases, the magnetic field decreases rapidly in certain directions, which demonstrates that not all magnetic fields get amplified in the cosmic string wake. We find that the amplification or decay depends on the length scale of the magnetic field perturbation. A detailed analysis reveals that the decay of the magnetic field is due to the breakdown of Alfven’s theorem at very small lengthscales. In the numerical set up, we use a exponentially decreasing magnetic field to illustrate our findings.
Our detailed study also shows that if the sheared magnetic field perturbation in the cosmic string wake is aligned to the flow of the plasma past the string, then the magnetic field is always amplified for whatever perturbation is given. It is only when the magnetic field perturbation is perpendicular to the direction of the flow, the decay lengthscale of the perturbed magnetic field determines whether the field gets amplified or whether it decays. This is because the frozen in magnetic field is affected more by the perpendicular component of the velocity of the fluid rather than the parallel component. We have presented results for a cosmic string moving in a high $\beta$ plasma as the wake formation in our numerical study is more pronounced for the high $\beta$ plasmas. For the numerical part, we have used the OpenMHD code, an open source code, which is available from GitHub.
@@
In section II, we discuss the spacetime around a cosmic string and discuss how we generate vorticity in the cosmic string wake structure. In section III, we discuss the amplification of the magnetic field due to the deformation of the magnetic flux lines. In Section IV, we show the decay of the magnetic field in presence of a sheared magnetic field. We also show that in the presence of a sheared magnetic field, the amplification or decay depends on the orientation and the lengthscale of the perturbative field. Section V we summarize the work and discuss why it is important to understand the evolution of magnetic fields in the cosmic string wakes for current observational signatures.
2 Vorticity in a cosmic string wake
@@
The geodesic equations around a cosmic string metric have been studied in detail in the literature abhisek . It has been shown that particles can have closed or open orbits around an Abelian Higgs string based on their total angular momentum and energy. In general, particles moving past a cosmic string suffer a change in their velocities close to the cosmic string. This leads to the wake formation behind moving cosmic strings. The metric being symmetric around the deficit angle of the string, the wake structure is also symmetric around the deficit angle of the string.
@@
The metric of a cosmic string is given by vilenkin ; deruelle ,
$ds^{2}=N^{2}(\rho)dt^{2}-d\rho^{2}-L^{2}(\rho)d\theta^{2}-N^{2}(\rho)dz^{2}$ (1)
The factors $L(\rho)$ and $N(\rho)$ are determined by the boundary conditions. They are related to the values of the fields at a distance $\rho$. The angle $\theta$ varies from 00 to $(1-4G\mu)\times 2\pi$. The missing angle $\Delta\theta=8\pi G\mu$ is referred to as the deficit angle of the cosmic string.
Apart from the presence of the deficit angle, the metric of the cosmic string is locally flat. So, the ideal magnetohydrodynamic (MHD) equations remain the same. They are given by mkverma ,
@@
The only difference comes as the particles move past the string. The conical nature of the metric gives a velocity perturbation to the particle towards the cosmic string. This velocity perturbation is proportional to the deficit angle $\Delta\theta$ of the string. It is given by $\delta v\approx\delta\theta v_{s}\gamma_{s}$, where, $v_{s}$ is the cosmic string velocity and $\gamma_{s}$ is the Lorentz gamma factor corresponding to the moving cosmic string. But this perturbation does not generate a local vorticity. The cosmic string metric is conical, which means that it is locally flat. From the reference frame of the particle, the path of the particle is always a straight line unless it is acted upon by some external force. Since the particles always move in a straight line, no local vorticity can be generated by the particles.
From an external observer’s reference frame however, the trajectories of the particles are bent. This gives rise to the overdensity of the fluid behind a moving cosmic string. However, massive particles which constitute the plasma of the early universe do have an angular momentum if the cosmic strings are Abelian Higgs strings hartmann . It has been shown that massive test particles can have a bound orbit or an unbound circular orbit around a Abelian Higgs cosmic string. It is therefore possible that the trajectory of the particle may not be a straight line in the reference frame of the particle. In this work, we do not specify the nature of the cosmic string, but we assume that the velocity of the particle moving in the plasma is affected by the collisions with other plasma particles. We consider a hydrodynamic approach where the velocity space of the number density of the particles in a plasma is treated as a distribution function. We generate a local vorticity in the wake behind the string by giving a perturbation to the velocity in both the $x$ and $y$ directions. These are the two directions in which the plane of the wake is considered. We consider these perturbations to the particle velocity to be of the order of the $\delta v$.
The perturbation to the velocity in both the directions may be due to various reasons. There are density inhomogeneities generated at several lengthscales in the wake of the cosmic string. The different species of particles in the wake cluster at different lengthscales. The collision between these particles may generate velocity perturbations in the plasma. Particle clustering non-uniformly close to an Abelian Higgs string has been discussed in previous hydrodynamic studies of wake structures sornborger , so small perturbations to the velocity of the particles is quite possible. This is why we give a small initial perturbation to the velocity in both the planar directions of the wake. This leads to the generation of a local vorticity in the plasma. This is shown if fig 1.
Vorticity in magnetohydrodynamic wakes are more common than in hydrodynamic wakes. This is because of the presence of the magnetic field. If we consider the trajectory of a charged particle, such as an electron in the magnetized plasma in a flat space time, the equation of motion is given by,
$m_{e}[\frac{\partial\mathbf{v_{e}}}{\partial t}+(\mathbf{v_{e}}.\nabla)\mathbf{v_{e}}]=\frac{-\nabla P_{e}}{n_{e}}+e_{e}\mathbf{E}+\frac{q}{c}\mathbf{v_{e}}\times\mathbf{B}$ (5)
Here the subscript ${e}$ denotes the electron as the charged particle but such an equation can be written down for any species of charged particles. Here $m_{e}$ denotes the mass of the particle and $P_{e}$, the hydrodynamic pressure. $n_{e}$ is the number density of the charged particles. In the presence of a magnetic field therefore, it is possible to define a generalized vorticity for the fluid. Charged particles move in rotational motion in a magnetic field and the generalized vorticity is defined by mahajan ,
$\boldsymbol{\omega_{G}}=\triangledown\times\mathbf{v_{e}}+\frac{q\mathbf{B}}{c}$ (6)
This equation is valid as long as the density is a function of pressure. In the case of the ion being the charged particle, similar equations can be written down. To obtain the one fluid MHD equations, some approximations have to be made. The mass of the electron is considered to be negligible compared to the mass of the ions. The plasma is assumed to be quasi neutral, $|n_{e}-n_{i}|<<n_{e}$. It is also assumed that the dispacement current from the Maxwell’s equation is negligible. Based on these assumptions, the ideal one fluid MHD equations can be used to study the evolution of the magnetic field.
As is well known, for an ideal fluid, the vorticity generated in the fluid will always be conserved. So once we generate the vorticity due to the velocity perturbations in the plasma it remains constant. After generating the vorticity, we use the one fluid model of the MHD equations to evolve the different fields (pressure, magnetic field etc). The vorticity remains constant throughout the evolution of the wake structure and therefore we do not study it’s evolution with time. Moreover, since the magnetic field lines are frozen in the plasma, the evolution of the equations of motion of the charged particles will be related to both the magnetic pressure as well as the hydrodynamic pressure in the fluid. We show later on why “coarse-graining” the fields and the smallest lengthscale are important for the growth or decay of the magnetic field in the magnetized plasma.
2.1 Numerical Simulation of the vorticity
@@
We use the publicly available OpenMHD code openmhd for our simulations. We use the basic code and use open boundary conditions in the direction of the fluid flow. In the basic code, the MHD equations along with the equation of state, relating the hydrodynamic pressure and the density, $P_{H}=f(\rho)$ are solved using a perturbative method. For the case of magnetogydrodynamics, the total pressure ($P$) consists of the hydrodynamic pressure and the pressure due to the magnetic field mkverma . The fields ($\mathbf{v}$, $\mathbf{B}$, $P$, $\rho$) mentioned previously in the MHD equations can be split into a background field and a perturbative field as, $\mathbf{v}=\mathbf{v_{0}}+\mathbf{v_{1}}$, $\mathbf{B}=\mathbf{B_{0}}+\mathbf{B_{1}}$, $P=P_{0}+P_{1}$ and $\rho=\rho_{0}+\rho_{1}$. Here $(\mathbf{v_{0}}$, $\mathbf{B_{0}}$, $P_{0}$, $\rho_{0})$ are the background fields, and $(\mathbf{v_{1}}$, $\mathbf{B_{1}}$, $P_{1}$, $\rho_{1})$ are the perturbative fields. A Galilean transformation is used to shift to the background flow frame, thus, $\mathbf{v_{0}}=0$, $\mathbf{v_{1}}=\mathbf{v}$ which means that the velocity field $\mathbf{v}$ is the perturbative field. These are then substituted in the MHD equations mentioned previously, and only the first order terms are retained, assuming the perturbations to be small. The linearized equations are given by,
@@
The basic code is available in one and two dimensions. We have used the 2D code and modified it by giving a velocity perturbation in both the directions of the plane of the wake to the particles as they cross the cosmic string. The fluid flow is taken to be along the $x$-axis. In the $y$-direction, we have used fixed boundary conditions. The cosmic string is aligned with the $z$-axis, it is identified in the flow by the position at which an initial velocity perturbation is given to the plasma particles. Thus, at the initial time, plasma particles crossing a certain position ($x$ coordinate) are all given a velocity perturbation both in the $x$-and $y$-direction; $v_{x}=-v_{0}Cos\theta$ and $v_{y}=-v_{0}Sin\theta$. All other particles move with velocity $-v_{0}$. This is in the reference frame of the cosmic string. As the particles flow, after some time, the vortex structure is formed. Fig.1 shows a streamline plot of the vortex structure generated in a cosmic string wake.
Figure 1: Generation of a pair of vortex in the cosmic string wake.
@@
For a better understanding of the simulation, a larger deficit angle is taken so that the vorticity is clearly visible. A larger deficit angle will result in a bigger dimension of the wake. We prefer to keep the order of the velcocity perturbations that we give to the particles of the same order as the velocity perturbation due to the cosmic string metric. As mentioned before, the deficit angle for an actual cosmic string is very small, in dimensionless parameters the perturbation would be of the order of the deficit angle with a value of $10^{-5}$ gangui . If we give a perturbation of the same order to the velocity along both the dimensions of the wake, the lengthscale of the vortex will be of the order of $10^{-5}Mpc$ or smaller. In our simulations, the net vorticity over a certain region is calculated and in the absence of a magnetic field, we find that the vorticity is conserved throughout the flow. This is an important check, as we have mentioned that we do not consider any dissipative forces in the system, so the energy conservation will also mean the conservation of vorticity in the flow.
2.2 Magnetic fields in rotational flows

### v3

@@ Abstract
We do a detailed study on the evolution of a magnetic field in a magnetized plasma within the spacetime of a moving cosmic string using analytical and numerical methods. The conical spacetime around the cosmic string causes the frozen-in magnetic field to deform due to the fluid flow. We find that the overdensity in the wake region amplifies the magnetic field. This amplification depends on the direction and the lengthscale of the magnetic perturbations. Alfven’s theorem of flux conservation explains this result. However, our study also shows that the magnetic field can decay depending on the perturbation lengthscale, due to the breakdown of Alfven’s theorem at a certain lengthscale. We have identified this lengthscale as the gyroradius of the charged particles in the plasma. Our findings are significant for understanding magnetic reconnection in cosmic string wakes.
magnetohydrodynamics, cosmic string wake
I Introduction
@@
Cosmic strings are topological defects that are formed from the symmetry-breaking phase transitions in the early universe. An interesting part of these defects is the nature of the spacetime around the string. It is known that the cosmic string metric is globally conical and has a deficit angle. Due to the nature of the metric, the geodesics of particles close to the cosmic string move inwards towards the cosmic string. This gives rise to a wake structure behind a moving cosmic string. This wake structure is well studied for an un-magnetized plasma vishniac ; sornborger ; stebbins . It has been found that these wakes leave several signatures during the evolution of the universe. They can give rise to discontinuities in the Cosmic Microwave Background Radiation (CMBR) danos ; they may give rise to primordial magnetic fields vachaspati1 ; vachaspati2 ; sovan and they may also lead to density inhomogeneities which will affect the nucleosynthesis calculations layek1 ; layek ; sovan1 . Thus, understanding the structure of these wakes is important for identifying the signatures of cosmic strings in the current universe. Currently, due to the availability of a large amount of data and computational resources, there has been a renewed interest in the experimental signature of these exotic objects Planck . Some of these are from density inhomogeneities and some from Gravitational Waves (GW) sousa . There are also some signatures from synchrotron radiation and Gamma Ray Bursts (GRB) cyr . Density inhomogeneities are generated in the wakes of cosmic strings. For the generation of density inhomogeneities it is necessary to study the hydrodynamic wake structure. But for signals like synchrotron radiation and GRBs one must consider the magnetic field in the background plasma. The wake structure that has been studied in the literature is essentially the hydrodynamic wake structure. The presence of magnetic fields in the wake was generally attributed only to the superconducting cosmic strings rogozin . Recently, it has been shown that primordial magnetic fields can be generated in the wakes of Abelian Higgs strings sovan too. So, it is possible that wakes of non-superconducting strings will also have a magnetic field. Particles moving through these magnetic fields will radiate synchrotron radiation soumen2 . Such radiation can be detected by current detectors. We need to study the electron distribution, the magnetic field perturbation and many other parameters before we are able to obtain a distinct signature that points to a cosmic string. Our motivation in this study is to study in greater detail the flow of the magnetized fluid around a cosmic string and understand the important lengthscales associated with this flow. This is an important factor for the radiation emitted from magnetic reconnections in the cosmic string wake.
@@
In our previous studies, we have shown that multiple shocks can be generated in the cosmic string wake soumen . These shocks can collide and release a large amount of energy as radiation. Moreover, it has also been conjectured that the narrow wake structure of the cosmic strings can lead to magnetic reconnection dilip . The shocks in the wake are usually enhanced when there is magnetic reconnection and the radiation that is emitted can be of the order of a GRB DilipGRB . Lengthscales play an important role in magnetic reconnection. One of the requirements of magnetic reconnection is the breakdown of Alfven’s theorem in the magnetized plasma. This can happen at small lengthscales. We aim to identify the lengthscale at which magnetic reconnection can occur in cosmic string wakes. This lengthscale is important as it determines the energy released due to magnetic reconnection.
@@
The fluid flow around the cosmic string is driven by the conical metric of the cosmic string. In the pre-recombination era, the magnetic field lines in the early universe are usually frozen into the plasma. Studies have shown that when the plasma undergoes a spherical compression, the magnetic field frozen in the plasma is also affected king . Since the flow around the cosmic string results in the fluid density increasing significantly in the wake region, we do a detailed study of the evolution of frozen-in magnetic fields around a cosmic string.
@@
In the literature, the magnetic field in the wake region is generated due to the presence of density inhomogeneities in the plasma. These density inhomogeneities come from the clustering of different particles around the cosmic string. The clustering of these particles mostly depend on their masses abhisek . A long and narrow wake is formed. Previously we have shown in numerical simulations, that when the velocity of the cosmic string is low, the wake structure is diffuse soumen . A well developed wake structure is obtained for higher Mach numbers. For simplicity, an ideal fluid approximation is considered without any dissipative forces. To simulate the wake like structure in a locally flat geometry, the knowledge of the boosted metric of a cosmic string is used to initiate a velocity perturbation in the flow of particles as they pass by the cosmic string.
Though hydrodynamic wakes and shocks are well studied for a cosmic string, it is only in recent times that magnetohydrodynamic shocks have also been studied in the cosmic string wake. Magnetohydrodynamic shocks are difficult to study as the magnetic field lengthscales play a major role in the generation and evolution of the shock dahlburg . Among cosmic string wakes, MHD shocks have been discussed previously in ref. beresnyak . In recent times, more detailed simulations of such wakes have been carried out in ref. soumen . One of the crucial factor, that was not considered previously in these studies was the lengthscale of the perturbations that affect the magnetic and density fields. Lengthscales of both density inhomogeneities as well as magnetic fields affect structure formation and later evolution of large scale structures in the universe kim . For cosmic strings, this is particularly important as density perturbation depend on the small-scale structure of the strings wu as well as the nature of the particle motion around them abhisek . In this respect, our study clearly shows the effect of magnetic perturbations of different lengthscales on the evolution of the magnetic field. We find that there is a lengthscale beyond which the magnetic field is not amplified any more. This is the lengthscale at which Alfven’s theorem breaks down, Our analytical and numerical studies show that this lengthscale corresponds to the gyro radius of the charged particle in the magnetic field.
@@
In section II, we discuss the spacetime around a cosmic string and why a moving cosmic string will always generate a velocity perturbation on the particles moving past it. In section III, we discuss the amplification of the magnetic field due to the deformation of the magnetic flux lines. In Section IV, we show the decay of the magnetic field in presence of a sheared magnetic field. We also show that in the presence of a sheared magnetic field, the amplification or decay depends on the orientation and the lengthscale of the perturbative field. Section V we summarize the work and discuss why it is important to understand the evolution of magnetic fields in the cosmic string wakes for current observational signatures.
II The cosmic string metric and wake formation
@@
The geodesic equations around a cosmic string metric have been studied in detail in the literature abhisek ; hartmann . In general, particles moving past a cosmic string suffer a change in their velocities close to the cosmic string. This leads to the wake formation behind moving cosmic strings sornborger . The metric being symmetric around the deficit angle of the string, the wake structure is also symmetric around the deficit angle of the string. Here, we would like to briefly recapitulate the basics of the metric of a cosmic string and discuss the velocity perturbation on particles going past a long string.
@@
The simplest metric of a cosmic string in cylindrical coordinates is given by vilenkin ; deruelle ,
$ds^{2}=dt^{2}-d\rho^{2}-(1-4G\mu)^{2}\rho^{2}d\theta^{2}-dz^{2}$ (1)
Here, $\mu$ is the mass per unit length of the cosmic string and comes from the symmetry breaking scale at which the cosmic string was formed. A cosmic string always has a finite width. Let us denote the width of the cosmic string by $r_{w}$. Beyond the width of the cosmic string the exterior metric of the string can be written as,
$ds^{2}=dt^{2}-d\rho^{2}-\rho^{2}d\theta^{\prime 2}-dz^{2}$ (2)
where, the $\theta^{\prime}$ now goes from 0 to $2\pi(1-4G\mu)$. In the Cartesian coordinate system, this leads to the standard flat metric,
$ds^{2}=dt^{2}-dx^{2}-dy^{2}-dz^{2}$ (3)
with the definition of $x=\rho~Cos~\theta^{\prime}$ and $y=\rho~Sin~\theta^{\prime}$. This is just the Minkowski space time with a deficit angle. Thus, this means that the standard MHD equations in Minkowski space time can be used to understand the flow around a cosmic string provided the boundary conditions are set so that the angle $\theta^{\prime}$ is constrained to be within 0 to $2\pi(1-4G\mu)$.
However, there is a difference in the case of the moving cosmic string. This difference lies in the fact that a moving particle is subject to a velocity perturbation towards the cosmic string. It is given by $\delta v\approx\delta\theta v_{s}\gamma_{s}$, where, $v_{s}$ is the cosmic string velocity and $\gamma_{s}$ is the Lorentz factor corresponding to the moving cosmic string. This is because of the large mass per unit length of the string which is like a gravitational force and can be written as a metric perturbation obtained from the linearized Einstein’s equations for the cosmic string, details of this are provided in ref. vachaspati2 . A particle at rest with respect to a string is not acted upon by any gravitational force but if the string moves with a velocity $v_{s}$ nearby matter particles are subject to a boost. The exact expression of the velocity perturbation depends on the nature of the metric and therefore on the particular kind of cosmic string that we are working with. One can obtain the most general form of the boosted metric in the Cartesian coordinate as Maarten ,
$ds^{2}=dt^{2}-dx^{2}-dy^{2}-dz^{2}-4G\mu(2-4G\mu)\left(\frac{(ydx-xdy)^{2}}{(x^{2}+y^{2})}\right)$ (4)
In this case, since $4G\mu<<1$, the last term can be treated as a metric perturbation on the flat space time. However, as has been discussed in detail in ref. rezolla , this metric perturbation does have a role to play in changing the geodesics of the particles flowing around the cosmic string. It has been shown that for both long strings as well as wiggly cosmic strings this results in a velocity perturbation proportional to the deficit angle of the cosmic string vachaspati1 ; vachaspati2 .
The metric is written as $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, where $\eta_{\mu\nu}$ is the metric for the flat space time and $h_{\mu\nu}$ is the perturbation as $4G\mu<<1$. So, the geodesic equation of motion will only be affected by the metric perturbation in the $x$- and $y$-direction. The equations of motion with the perturbed metric is a rather complex equation which can be solved numerically. The nature of the geodesic equations in the $x$- and $y$-directions are,
$\frac{dv_{x}}{dt}=-4G\mu(2-4G\mu)(\frac{v_{x}^{2}}{c}f_{1}(x,y)+\frac{v_{x}v_{y}}{c}f_{2}(x,y)+\frac{v_{y}^{2}}{c}f_{3}(x,y))$ (5)
$\frac{dv_{y}}{dt}=-4G\mu(2-4G\mu)(\frac{v_{y}^{2}}{c}f_{1}(x,y)+\frac{v_{x}v_{y}}{c}f_{2}(x,y)+\frac{v_{x}^{2}}{c}f_{3}(x,y))$ (6)
Here, the functions $f_{1}$, $f_{2}$, $f_{3}$ are functions of the variables $x,y$ and the constants $G,\mu$. They are the functions involving squared and cubic terms and can only be solved numerically but they do indicate that there would be a velocity change on the plasma particles due to the motion of the cosmic string. While studying accretion of particles in cosmic string wakes, a novel method is used to get this velocity perturbation. If a particle has a velocity $\vec{v}$ and it’s trajectory has to be computed, it is given a Lorentz transformation to the cosmic string’s rest frame and the trajectory is computed in that frame of reference, it is then transformed back to obtain the change in velocity due to the moving string. In ref. bertschinger , the velocity change for a relativistic string moving along the $y$-direction has been found. Since the cosmic string in our case is moving along the x-axis, the velocity perturbation due to the moving string in our case would be,
$\Delta v_{x}=-4\pi\frac{G\mu\gamma_{s}}{c^{2}}(v_{s}-\frac{v_{s}^{2}v_{x}}{c^{2}})$ (7)
$\Delta v_{y}=-4\pi\frac{G\mu\gamma_{s}}{c^{2}}(v_{s}-v_{x}-\frac{v_{s}v_{y}^{2}}{c^{2}})$ (8)
Here, $v_{s}$ is the velocity of the string.
There are different methods to obtain the velocity perturbations due to moving cosmic strings. Since the velocity perturbation depends on the exact metric of the cosmic string, it will be numerically different for different kinds of cosmic string. But the important point that we see in the two calculations is that the velocity perturbation is always proportional to the symmetry breaking scale of the cosmic string i.e. $\Delta v\propto 4\pi G\mu$. So, particles moving past a cosmic string will move in a locally flat space time but will be subjected to a velocity perturbation as they pass by the cosmic string. The metric though flat will have a condition on the angle $\theta$ that the total angle in the space should sum upto $2\pi(1-4G\mu)$ only. Any simulation of cosmic string wakes should satisfy these two conditions.
Wake formation from cosmic string has been studied in hydrodynamics previously sornborger and it has been shown that particles with different masses would cluster differently in the cosmic string wakes. In recent times, motion of long cosmic strings have been studied in a magnetized plasma using the magnetohydrodynamic equations in locally flat space time but with a deficit angle soumen . Loops of magnetic reconnection has been observed in the simulation. The possibility of having magnetic reconnection in cosmic string wakes have also been shown in a separate work dilip . Since magnetic reconnection can lead to various observable signatures, it is important to understand the lengthscales involved in the evolution of the magnetic fields in magnetized cosmic string wakes.
II.1 Numerical Simulation of flow around a cosmic string
For the numerical simulation of fluid flow around a cosmic string, we would like to point out that generally fluid particles in an ideal fluid follow the geodesic of the underlying metric, only in this case the differential equations denoting the flow of the fluid are not solved for individual particles but for a volume of particles as a whole. So it is the fluid density that is used in the equation and not the individual mass and trajectory of each particle. There has been some formulation of the mapping of the geodesics into the hydrodynamic formulation in the literature. The fluid flow around the cosmic string is beyond the core of the cosmic string so in this region $\rho>>r_{w}$. As the external metric of the cosmic string, determines the flow of the particles in this region, in Cartesian coordinates, it is given by the metric in Eq.(3), the MHD equations in these coordinates would be given by mkverma ,
@@ II.1 Numerical Simulation of flow around a cosmic string
The important thing to keep in mind is that the angle $\theta$ in the xyx-y plane is defined only between 0 to $2\pi(1-4G\mu)$.
@@
We use the publicly available OpenMHD code openmhd for our simulations. We use the basic code and use open boundary conditions in the direction of the fluid flow. In the basic code, the MHD equations along with the equation of state, relating the hydrodynamic pressure and the density, $P_{H}=f(\rho)$ are solved using a perturbative method. For the case of magnetohydrodynamics, the total pressure ($P$) consists of the hydrodynamic pressure and the pressure due to the magnetic field mkverma . The fields ($\mathbf{v}$, $\mathbf{B}$, $P$, $\rho$) mentioned previously in the MHD equations can be split into a background field and a perturbative field as, $\mathbf{v}=\mathbf{v_{0}}+\mathbf{v_{1}}$, $\mathbf{B}=\mathbf{B_{0}}+\mathbf{B_{1}}$, $P=P_{0}+P_{1}$ and $\rho=\rho_{0}+\rho_{1}$. Here, $(\mathbf{v_{0}}$, $\mathbf{B_{0}}$, $P_{0}$, $\rho_{0})$ are the background fields, and $(\mathbf{v_{1}}$, $\mathbf{B_{1}}$, $P_{1}$, $\rho_{1})$ are the perturbative fields. A Galilean transformation is used to shift to the background flow frame, thus, $\mathbf{v_{0}}=0$, $\mathbf{v_{1}}=\mathbf{v}$ which means that the velocity field $\mathbf{v}$ is the perturbative field. These are then substituted in the MHD equations mentioned previously, and only the first order terms are retained, assuming the perturbations to be small. The linearized equations are given by,
@@
The basic code is available in one and two dimensions. We have used the 2D code and modified it by giving a velocity perturbation in the plane of the wake to the particles as they cross the cosmic string. The fluid flow is taken to be along the $x$-axis. In the $y$-direction, we have used fixed boundary conditions. The cosmic string is aligned with the $z$-axis, it is identified in the flow by the position at which an initial velocity perturbation is given to the plasma particles. Thus, plasma particles crossing a certain position are all given a velocity perturbation in the $x$-and $y$-direction. The velocity perturbation is given by Eq.(7) and Eq.(8). But this velocity perturbation is for each individual particle moving along a geodesic. Assuming that the velocity of the particles are less than the velocity of the cosmic string which moves close to the speed of light, we consider the average velocity perturbation to be of the order of $4\pi G\mu v_{r}$, where $v_{r}$ is the relative velocity between the particle and the cosmic string. This average velocity is denoted by $v_{0}$. To see that the constraint of the angle is maintained in the xyx-y plane, we take $v_{x}=-v_{0}Cos\theta$ and $v_{y}=-v_{0}Sin\theta$. The angle $\theta$ is proportional to the deficit angle. It can neither be zero nor $2\pi$ as we know from the discussion in the previous section that the total angular deviation is constrained by the deficit angle of the metric. For all other volume elements which are away from the cosmic string, the velocity perturbation is zero. Thus the velocity perturbation is only given when the plasma moves past the cosmic string.
@@
Initially, for a better understanding of the simulation, a larger deficit angle was taken so that the results are clearly visible. A larger deficit angle will result in a bigger dimension of the wake. However, we prefer to keep the order of the velocity perturbations that we give to the particles of the same order as the velocity perturbation due to the cosmic string metric. As mentioned before, the deficit angle for an actual cosmic string is very small, in dimensionless parameters the perturbation would be of the order of the deficit angle with a value of $10^{-5}$ gangui . If we give a perturbation of the same order to the velocity along both the dimensions of the wake, the lengthscale of the wake region will be of the order of $10^{-5}Mpc$ or smaller.
II.2 Magnetic fields in rotational flows